# Multidimensional Dickman distribution and operator selfdecomposability

### v1

@@ 2 One-dimensional Dickman distribution
The one-dimensional Dickman distribution with parameter $\theta>0$ can be defined as a distribution of a random variable $X_{D}$ satisfying the equation
@@
where the symbol $\overset{d}{=}$” denotes the equality in distribution, $X^{\prime}_{D}\overset{d}{=}X_{D}$, $U$ has the uniform distribution on $[0,1]$ and is independent of $X^{\prime}_{D}$, see e.g. [PW04]. Equivalently, the random variable $X_{D}$ is the fixed point (in a distributional sense) of the random affine transformation $\cdot\mapsto U^{1/\theta}(1+\cdot)$, where $U$ is as above. We will further denote $GD_{\theta}$ the Dickman distribution with parameter $\theta>0$.
@@
The density $f_{\theta}(x)$ of the $GD_{\theta}$ distribution has been an object of independent studies, and we recall here a few results related to it. The asymptotic behaviour of the density $f_{\theta}(x)$ for $\theta=1$ has been established in [DE 51b] and for general $\theta$ in [VER72, Lem. 4.7.9]. It was also proven in [VER72, Thm. 4.7.7] that the density $f_{\theta}(x)$ can be expressed explicitly as follows
@@ 3 Dickman distributions over $\mathbb{R}^{d}$
Let $\|Q\|\coloneq\sup_{|\mathbf{x}|\leq 1}|Q\mathbf{x}|$ denote the operator norm of $Q\in\mathscr{M}$. For any measure $\vartheta$ on $\mathbb{R}^{d}$ and any invertible operator $T:\mathbb{R}^{d}\mapsto\mathbb{R}^{d}$ denote $T\vartheta$ the measure given by $T\vartheta(A)\coloneq\vartheta(T^{-1}(A))$, $A\in\mathscr{B}(\mathbb{R}^{d})$.
@@
and let $\mathscr{S}\subset\mathscr{P}(\mathbb{R}^{d})$ be the set of all probability distributions concentrated on $d$-dimensional sphere, i.e. $\sigma\in\mathscr{S}$ whenever $\sigma(\mathbb{S}^{d-1})=1$. Note that $\mathscr{S}\subset H$. For a random vector $\mathbf{X}\in\mathbb{R}^{d}$ let $\mathscr{L}(\mathbf{X})\in\mathscr{P}(\mathbb{R}^{d})$ denote its probability distribution.
@@
Next, we generalise this definition to a wider class of distributions by substituting the scalar function $u^{1/\theta},\theta>0$ with a matrix exponential $u^{Q},Q\in\mathscr{M}_{+}$ and a probability measure $\sigma\in\mathscr{S}$ with a probability measure $\nu\in H$.
@@
Moreover, for any $\mathscr{Q}\subseteq\mathscr{M}_{+}$ and $\mathscr{H}\subseteq H$ let $\mathscr{D}(\mathscr{Q},\mathscr{H})$ be a set of distributions defined as
@@
The existence of a weak limit of $\mathbf{Y}_{n}$ as $n\to\infty$, independent of the initial condition $\mathbf{Y}_{0}$, under some mild conditions on $M_{n}$ and $\bm{\xi}_{n}$ was established in the celebrated work of Kesten [KES73]. In particular, Kesten [KES73, Thm. 6] showed that the limit exists if $\infty<\mathbb{E}\log\|M_{1}\|<0$ and $0<\mathbb{E}|\bm{\xi}_{1}|^{\kappa}<\infty$ for some $\kappa>0$. However, as shown in [VER72] for $d=1$, the last condition can be weakened to $\mathbb{E}\log^{+}|\xi_{1}|<\infty$ under the assumption that the first condition holds, and this proof translates easily for $d>1$. Hereafter, we define
@@
(i) First note that the function $f(t)=t\wedge 1$ is non-decreasing. Note also that for any $Q\in\mathscr{M}_{+}$ there exist $K>0$ and $c_{1}>0$ such that $|s^{Q}\mathbf{x}|\leq c_{1}s^{K}|\mathbf{x}|$ for all $s>0$ and $\mathbf{x}\in\mathbb{R}^{d}$ [MS01, Prop. 2.2.11], see also the proof of [URB72, Lem. 6.1.]. We might as well choose $c_{1}>1$, thus
$\int\limits_{0}^{1}(|s^{Q}\mathbf{x}|\wedge 1)\frac{ds}{s}$ $\leq c_{1}\int\limits_{0}^{1}(s^{K}|\mathbf{x}|\wedge 1)\frac{ds}{s}=\frac{c_{1}}{K}\int\limits_{0}^{1}(s|\mathbf{x}|\wedge 1)\frac{ds}{s}$
$=\frac{c_{1}}{K}\int\limits_{0}^{|\mathbf{x}|}(s\wedge 1)\frac{ds}{s}=\frac{c_{1}}{K}\left(1\wedge|\mathbf{x}|-\log\left(1\wedge|\mathbf{x}|^{-1}\right)\right),$
@@
Letting $r=0$ we can deduce from $\psi(\mathbf{0})=1$ that $g(\bm{\omega})=0$ for all $\bm{\omega}\in\mathbb{R}^{d}$.
@@
Remarkably, the distributions with Lévy measure as in (12) have previously appeared in [JUR82].
@@
    Assume that $\nu\in H$ is such that $\int\limits_{\mathbb{R}^{d}}|\mathbf{x}|\nu(\mathbf{x})<\infty$. Then the mean vector $\mathbf{m}_{Q,\nu}$ of $\mathscr{D}(Q,\nu)$ exists for any $Q\in\mathscr{M}_{+}$ and is given as
@@
    Assume that $\nu\in H$ such that $\int\limits_{\mathbb{R}^{d}}|\mathbf{x}|^{2}\nu(\mathbf{x})<\infty$. Then the covariance matrix $\mathbb{C}_{Q,\nu}$ of $\mathscr{D}(Q,\nu)$ exists for any $Q\in\mathscr{M}_{+}$ and can be given by
@@
Both formulas (13) and (14) follow easily by differentiation of the characteristic function (11). Note that the finiteness of the integrals in (13) under the assumptions of (i) follows again by the estimate
$|s^{Q}\mathbf{x}|\leq{c_{1}}s^{K}|\mathbf{x}|,\;s>0,\mathbf{x}\in\mathbb{R}^{d}$
with some $c_{1},K>0$. Consequently, we obtain
$|\mathbf{m}|\leq\frac{c_{1}}{K}\int\limits_{\mathbb{R}^{d}}|\mathbf{x}|\nu(d\mathbf{x})<\infty.$
@@
$\begin{gathered}\mathbf{m}_{Q,\nu}=Q^{-1}\int\limits_{\mathbb{R}^{d}}\mathbf{x}\nu(d\mathbf{x})\;\text{ and }\;\mathbb{C}_{Q,\nu}=\mathscr{A}_{Q}^{-1}\left(\int\limits_{\mathbb{R}^{d}}\mathbf{x}\mathbf{x}^{T}\nu(d\mathbf{x})\right),\end{gathered}$
@@
    For any fixed $Q\in\mathscr{M}_{+}$ the class $\mathscr{D}(\frac{1}{c}Q,H)$, where $c$ runs over $(0,\infty)$, is closed under finite convolutions.
@@
(iii) Note first that for fixed $Q\in\mathscr{M}_{+}$, $c>0$ and $\mathscr{D}=\mathscr{D}(\frac{1}{c}Q,\nu)$, from (11) by a change of variables we obtain
$\log\widehat{\mathscr{D}}(\mathbf{z})=\int\limits_{\mathbb{R}^{d}}\int\limits_{0}^{1}(e^{i\mathbf{z}\cdot s^{\frac{1}{c}Q}\mathbf{x}}-1)\nu(d\mathbf{x})\frac{ds}{s}=c\int\limits_{\mathbb{R}^{d}}\int\limits_{0}^{1}(e^{i\mathbf{z}\cdot s^{Q}\mathbf{x}}-1)\nu(d\mathbf{x})\frac{ds}{s}.$
Let $Q\in\mathscr{M}_{+}$, $c_{1},c_{2}>0$ and $\nu_{1},\nu_{2}\in H$ and consider the independent distributions $\mathscr{D}_{1}=\mathscr{D}(\frac{1}{c_{1}}Q,\nu_{1})$, $\mathscr{D}_{2}=\mathscr{D}(\frac{1}{c_{2}}Q,\nu_{2})$. If we denote $\mathscr{D}=\mathscr{D}_{1}*\mathscr{D}_{2}$, then
$\log\widehat{\mathscr{D}}(\mathbf{z})$ $=\int\limits_{\mathbb{R}^{d}}\int\limits_{0}^{1}(e^{i\mathbf{z}\cdot s^{Q}\mathbf{x}}-1)(c_{1}\nu_{1}+c_{2}\nu_{2})(d\mathbf{x})\frac{ds}{s}$
$=(c_{1}+c_{2})\int\limits_{\mathbb{R}^{d}}\int\limits_{0}^{1}(e^{i\mathbf{z}\cdot s^{Q}\mathbf{x}}-1)\frac{(c_{1}\nu_{1}+c_{2}\nu_{2})(d\mathbf{x})}{c_{1}+c_{2}}\frac{ds}{s},$
and it is easy to see that $\mathscr{D}=\mathscr{D}(\frac{1}{c_{1}+c_{2}}Q,\frac{c_{1}\nu_{1}+c_{2}\nu_{2}}{c_{1}+c_{2}})$. ∎
@@
Let us fix $\Lambda\in\mathscr{M}_{+}$ and assume that $\Lambda$ is diagonal, that is $\Lambda=diag(a_{1},\ldots,a_{d})$ for $a_{i}>0,i=1,\ldots,d$. We will also assume that $\Lambda\notin\mathbb{I}$. Let $\mathscr{L}(\mathbf{X})=\mathscr{D}(\Lambda,\nu)$ for some $\nu\in H$. Since in this case it holds for $t>0$ that $t^{\Lambda}=diag(s^{a_{1}},\ldots,s^{a_{d}})$ it is clear that the components $X_{i}$ of the vector $\mathbf{X}$ satisfy the distributional equation
@@
$C_{Q,\nu}]_{i,j}=\frac{1}{a_{i}+a_{j}}\int\limits_{\mathbb{R}^{d}}x_{i}x_{j}\nu(d\mathbf{x}),\;i,j=1\ldots,d$
whenever $\int\limits_{\mathbb{R}^{d}}|\mathbf{x}|^{2}\nu(\mathbf{x})<\infty$. Notice also that in the simplest case $\nu=\delta_{\mathbf{w}}$, for some $\mathbf{w}=(w_{1},\ldots,w_{d})\in\mathbb{R}^{d}_{+}$, the components $w_{i}X_{i}$ are distributed according to $GD_{1/a_{i}}$, $i=1,\ldots,d$.
@@
Consider $\mathscr{D}=\mathscr{D}(\frac{1}{\theta}I,\nu)$, $\theta>0$ and assume $\nu\in\mathscr{S}$. As was noted before, the random vector $\mathbf{X}$ with $\mathscr{L}(\mathbf{X})=\mathscr{D}$ has the multidimensional Dickman distribution as defined in [BM20]. The properties of this distribution have been studied in [GZ24b], see also [GZ24a]. In particular, it was established that if the measure $\nu$ is finitely supported, then $\mathbf{X}$ can be expressed using independent Dickman random variables. Indeed, assume that $\nu$ is of the form
@@
In what follows, we focus on obtaining relations for the probability density function of the operator Dickman distributions. Notice first that if the distribution $\mathscr{D}(Q,\nu)$, $Q\in\mathscr{M}_{+}$, $\nu\in H$, over $\mathbb{R}^{d}$ is full, then it is absolutely continuous with respect to the Lebesgue measure on $\mathbb{R}^{d}$ [YAM83]. We also refer to [KEN99, Prop. 28.1] for sufficient conditions ensuring differentiability of the density for the distributions in the set $\mathscr{D}(\mathbb{I},\nu)$. Nevertheless, an explicit expression for the density is generally difficult to obtain. In this work, the explicit formula for the density is obtained for a specific case of the multidimensional Dickman distribution $\mathscr{D}\in\mathscr{D}(\mathbb{I},\mathscr{S})$ in Theorem 3.15, whereas Proposition 3.16 treats a more general case of $\mathscr{D}(\mathbb{I},H)$. Note that we utilise some elements of the theory of generalised functions operating on the Schwartz space of rapidly decreasing functions; for more details on this topic, see the classic book [GS69] and also [MIT13].
@@
We further denote $J_{a}(\cdot),a>-\frac{1}{2}$ the Bessel function of first kind, $Leb$ the Lebesgue measure on $\mathbb{S}^{d-1}$, $\zeta_{d}=(Leb(\mathbb{S}^{d-1}))^{-1}={\Gamma(d/2)}\cdot(2\pi)^{-d/2}$ the reciprocal surface area of $\mathbb{S}^{d-1}$ and $Y_{d}(s)=(s/2)^{1-\frac{d}{2}}\Gamma(d/2)J_{d/2-1}(s)$, $s>0$, $d>1$, the spherical Bessel function, see e.g. [PIN93]. Moreover, for $d=1$, we let $Y_{1}(s)=\cos(s),s>0$.
@@
Denote now $\tilde{F}_{k}(s)=s^{-d}\mathds{1}\{1<s<k\}$, $k\geq 2$ and $\tilde{F}(s)=s^{-d}\mathds{1}\{s>1\}$. Let also $F_{k}(\mathbf{x})=\tilde{F}_{k}(|\mathbf{x}|)$ and $F(\mathbf{x})=\tilde{F}(|\mathbf{x}|)$. Note that $\{F_{k},k\geq 2\}$ is a sequence of functions such that
@@
$\sum_{k=1}^{\infty}z_{k}\frac{\partial\hat{f}(\mathbf{z})}{\partial z_{k}}=\theta\hat{f}(\mathbf{z})\hat{\nu}(\mathbf{z})-\theta\hat{f}(\mathbf{z}).$

### v2

@@ 2 One-dimensional Dickman distribution
The one-dimensional Dickman distribution with parameter $\theta>0$ can be defined as the distribution of a random variable $X_{D}$ satisfying the equation
@@
where the symbol $\overset{d}{=}$” denotes the equality in distribution, $X^{\prime}_{D}\overset{d}{=}X_{D}$, $U$ has the uniform distribution on $[0,1]$ and is independent of $X^{\prime}_{D}$, see e.g. [PW04]. Equivalently, the random variable $X_{D}$ is the fixed point (in a distributional sense) of the random affine transformation $\cdot\mapsto U^{1/\theta}(1+\cdot)$, where $U$ is as above. We will further denote $GD_{\theta}$, the Dickman distribution with parameter $\theta>0$.
@@
The density $f_{\theta}(x)$ of the $GD_{\theta}$ distribution has been an object of independent studies, and we recall here a few results related to it. The asymptotic behaviour of the density $f_{\theta}(x)$ for $\theta=1$ has been established in [DE 51b] and for general $\theta>0$ in [VER72, Lem. 4.7.9]. It was also proven in [VER72, Thm. 4.7.7] that the density $f_{\theta}(x)$ can be expressed explicitly as follows
@@ 3 Dickman distributions over $\mathbb{R}^{d}$
Let $\|Q\|\coloneq\sup_{|\mathbf{x}|\leq 1}|Q\mathbf{x}|$ denote the operator norm of $Q\in\mathscr{M}$. For any measure $\vartheta$ on $\mathbb{R}^{d}$ and any measurable map $T:\mathbb{R}^{d}\mapsto\mathbb{R}^{d}$, denote $T\vartheta$ the measure given by $T\vartheta(A)\coloneq\vartheta(T^{-1}(A))$, $A\in\mathscr{B}(\mathbb{R}^{d})$.
@@
and let $\mathscr{S}\subset\mathscr{P}(\mathbb{R}^{d})$ be the set of all probability distributions concentrated on $d$-dimensional sphere, i.e. $\sigma\in\mathscr{S}$ whenever $\sigma(\mathbb{S}^{d-1})=1$. Note that $\mathscr{S}\subset H$. For a random vector $\mathbf{X}\in\mathbb{R}^{d}$, let $\mathscr{L}(\mathbf{X})\in\mathscr{P}(\mathbb{R}^{d})$ denote its probability distribution.
@@
Next, we generalise this definition to a wider class of distributions by substituting the scalar function $u^{1/\theta},\theta>0$, with a matrix exponential $u^{Q},Q\in\mathscr{M}_{+}$, and a probability measure $\sigma\in\mathscr{S}$ with a probability measure $\nu\in H$.
@@
Moreover, for any $\mathscr{Q}\subseteq\mathscr{M}_{+}$ and $\mathscr{H}\subseteq H$, let $\mathscr{D}(\mathscr{Q},\mathscr{H})$ be a set of distributions defined as
@@
The existence of a weak limit of $\mathbf{Y}_{n}$ as $n\to\infty$, independent of the initial condition $\mathbf{Y}_{0}$, under some mild conditions on $M_{n}$ and $\bm{\xi}_{n}$ was established in the celebrated work of Kesten [KES73]. In particular, Kesten [KES73, Thm. 6] showed that the limit exists if $\infty<\mathbb{E}\log\|M_{1}\|<0$ and $0<\mathbb{E}|\bm{\xi}_{1}|^{\kappa}<\infty$ for some $\kappa>0$. However, as shown in [VER79] for $d=1$, the last condition can be weakened to $\mathbb{E}\log^{+}|\xi_{1}|<\infty$ under some assumptions on $M_{1}$, and this proof translates easily for $d>1$ in our case, using (11) below. Hereafter, we define
@@
Note also that throughout the work we will make use of the following estimate. Let $Q\in\mathscr{M}_{+}$ be fixed, then there exist positive constants $c_{1},c_{2}$ and $k,K$ such that for all $0\leq s\leq 1$ it holds that
$c_{1}s^{K}\leq\|s^{Q}\|\leq c_{2}s^{k},$ (11)
see [MS01, Prop. 2.2.11] and the proof of [URB72, Lem. 6.1].
@@
(i) First note that the function $f(t)=t\wedge 1$ is non-decreasing. We can assume, without loss of generality, that $c_{2}>1$ in (11). Thus, we obtain
$\int\limits_{0}^{1}(|s^{Q}\mathbf{x}|\wedge 1)\frac{ds}{s}$ $\leq c_{2}\int\limits_{0}^{1}(s^{k}|\mathbf{x}|\wedge 1)\frac{ds}{s}=\frac{c_{2}}{k}\int\limits_{0}^{1}(s|\mathbf{x}|\wedge 1)\frac{ds}{s}$
$=\frac{c_{2}}{k}\int\limits_{0}^{|\mathbf{x}|}(s\wedge 1)\frac{ds}{s}=\frac{c_{2}}{k}\left(1\wedge|\mathbf{x}|-\log\left(1\wedge|\mathbf{x}|^{-1}\right)\right),$
@@
Letting $r=0$, we can deduce from $\psi(\mathbf{0})=1$ that $g(\bm{\omega})=0$ for all $\bm{\omega}\in\mathbb{R}^{d}$.
@@
Remarkably, the distributions with Lévy measure as in (13) previously appeared in [JUR82].
@@
    Assume that $\nu\in H$ is such that $\int\limits_{\mathbb{R}^{d}}|\mathbf{x}|\nu(d\mathbf{x})<\infty$. Then the mean vector $\mathbf{m}_{Q,\nu}$ of $\mathscr{D}(Q,\nu)$ exists for any $Q\in\mathscr{M}_{+}$ and can be given by
@@
    Assume that $\nu\in H$ such that $\int\limits_{\mathbb{R}^{d}}|\mathbf{x}|^{2}\nu(d\mathbf{x})<\infty$. Then the covariance matrix $\mathbb{C}_{Q,\nu}$ of $\mathscr{D}(Q,\nu)$ exists for any $Q\in\mathscr{M}_{+}$ and can be given by
@@
Both formulas (14) and (15) follow easily by differentiation of the characteristic function (12). Note that the finiteness of the integrals in (14) under the assumptions of (i) follows again from (11). Indeed, we can obtain
$|\mathbf{m}|\leq\frac{c_{2}}{k}\int\limits_{\mathbb{R}^{d}}|\mathbf{x}|\nu(d\mathbf{x})<\infty.$
@@
$\begin{gathered}\mathbf{m}_{Q,\nu}=Q^{-1}\int_{\mathbb{R}^{d}}\mathbf{x}\nu(d\mathbf{x})\;\text{ and }\;\mathbb{C}_{Q,\nu}=\mathscr{A}_{Q}^{-1}\left(\int_{\mathbb{R}^{d}}\mathbf{x}\mathbf{x}^{T}\nu(d\mathbf{x})\right),\end{gathered}$
@@
    For any fixed $Q\in\mathscr{M}_{+}$, the class $\mathscr{D}(w^{-1}Q,H)$, where $w$ runs over $(0,\infty)$, is closed under finite convolutions.
@@
(iii) Note first that for fixed $Q\in\mathscr{M}_{+}$, $w>0$ and $\mathscr{D}=\mathscr{D}(w^{-1}Q,\nu)$, from (12) by a change of variables we obtain
$\log\widehat{\mathscr{D}}(\mathbf{z})=\int\limits_{\mathbb{R}^{d}}\int\limits_{0}^{1}(e^{i\mathbf{z}\cdot s^{\frac{1}{w}Q}\mathbf{x}}-1)\nu(d\mathbf{x})\frac{ds}{s}=w\int\limits_{\mathbb{R}^{d}}\int\limits_{0}^{1}(e^{i\mathbf{z}\cdot s^{Q}\mathbf{x}}-1)\nu(d\mathbf{x})\frac{ds}{s}.$
Let $Q\in\mathscr{M}_{+}$, $w_{1},w_{2}>0$ and $\nu_{1},\nu_{2}\in H$ and consider the independent distributions $\mathscr{D}_{1}=\mathscr{D}(w_{1}^{-1}Q,\nu_{1})$, $\mathscr{D}_{2}=\mathscr{D}(w_{2}^{-1}Q,\nu_{2})$. If we denote $\mathscr{D}=\mathscr{D}_{1}*\mathscr{D}_{2}$, then
$\log\widehat{\mathscr{D}}(\mathbf{z})$ $=\int\limits_{\mathbb{R}^{d}}\int\limits_{0}^{1}(e^{i\mathbf{z}\cdot s^{Q}\mathbf{x}}-1)(w_{1}\nu_{1}+w_{2}\nu_{2})(d\mathbf{x})\frac{ds}{s}$
$=(w_{1}+w_{2})\int\limits_{\mathbb{R}^{d}}\int\limits_{0}^{1}(e^{i\mathbf{z}\cdot s^{Q}\mathbf{x}}-1)\frac{(w_{1}\nu_{1}+w_{2}\nu_{2})(d\mathbf{x})}{w_{1}+w_{2}}\frac{ds}{s},$
and it is easy to see that $\mathscr{D}=\mathscr{D}(\frac{1}{w_{1}+w_{2}}Q,\frac{w_{1}\nu_{1}+w_{2}\nu_{2}}{w_{1}+w_{2}})$. ∎
@@
Let us fix $\Lambda\in\mathscr{M}_{+}$ and assume that $\Lambda$ is diagonal, that is $\Lambda=diag(a_{1},\ldots,a_{d})$ for $a_{i}>0,i=1,\ldots,d$. We will also assume that $\Lambda\notin\mathbb{I}$. Let $\mathscr{L}(\mathbf{X})=\mathscr{D}(\Lambda,\nu)$ for some $\nu\in H$. Since in this case it holds for $s>0$ that $s^{\Lambda}=diag(s^{a_{1}},\ldots,s^{a_{d}})$, it is clear that the components $X_{i}$ of the vector $\mathbf{X}$ satisfy the distributional equation
@@
$\mathbf{C}_{Q,\nu}]_{i,j}=\frac{1}{a_{i}+a_{j}}\int\limits_{\mathbb{R}^{d}}x_{i}x_{j}\nu(d\mathbf{x}),\;i,j=1\ldots,d$
whenever $\int\limits_{\mathbb{R}^{d}}|\mathbf{x}|^{2}\nu(d\mathbf{x})<\infty$. Notice also that in the simplest case $\nu=\delta_{\mathbf{w}}$, for some $\mathbf{w}=(w_{1},\ldots,w_{d})\in\mathbb{R}^{d}_{+}$, the components $w_{i}X_{i}$ are distributed according to $GD_{1/a_{i}}$, $i=1,\ldots,d$.
@@
Consider $\mathscr{D}=\mathscr{D}(\frac{1}{\theta}I,\nu)$, $\theta>0$, and assume $\nu\in\mathscr{S}$. As was noted before, the random vector $\mathbf{X}$ with $\mathscr{L}(\mathbf{X})=\mathscr{D}$ has the multidimensional Dickman distribution as defined in [BM20]. The properties of this distribution have been studied in [GZ24b], see also [GZ24a]. In particular, it was established that if the measure $\nu$ is finitely supported, then $\mathbf{X}$ can be expressed using independent Dickman random variables. Indeed, assume that $\nu$ is of the form
@@
In what follows, we focus on obtaining relations for the probability density function of the operator Dickman distributions. Notice first that if the distribution $\mathscr{D}(Q,\nu)$, $Q\in\mathscr{M}_{+}$, $\nu\in H$, over $\mathbb{R}^{d}$ is full, then it is absolutely continuous with respect to the Lebesgue measure on $\mathbb{R}^{d}$ [YAM83]. We also refer to [KEN99, Prop. 28.1] for sufficient conditions ensuring differentiability of the density for the distributions in the set $\mathscr{D}(\mathbb{I},\nu)$. Nevertheless, an explicit expression for the density is generally difficult to obtain. In this work, the explicit formula for the density is obtained for a specific case of the multidimensional Dickman distribution $\mathscr{D}\in\mathscr{D}(\mathbb{I},\mathscr{S})$ in Theorem 3.15, whereas Proposition 3.16 treats a more general case of $\mathscr{D}(\mathbb{I},H)$. Note that we utilise some elements of the theory of generalised functions operating on the Schwartz space of rapidly decreasing functions; for more details on this topic, see the classical book [GS69] and also [MIT13].
@@
We further denote $J_{a}(\cdot),a>-\frac{1}{2}$, the Bessel function of first kind, $Leb$ the Lebesgue measure on $\mathbb{S}^{d-1}$, $\zeta_{d}=(Leb(\mathbb{S}^{d-1}))^{-1}={\Gamma(d/2)}\cdot(2\pi)^{-d/2}$ the reciprocal surface area of $\mathbb{S}^{d-1}$ and $Y_{d}(s)=(s/2)^{1-\frac{d}{2}}\Gamma(d/2)J_{d/2-1}(s)$, $s>0$, $d>1$, the spherical Bessel function, see e.g. [PIN93]. Moreover, for $d=1$, we let $Y_{1}(s)=\cos(s),s>0$.
@@
Denote now $\tilde{F}_{k}(s)=s^{-d}\mathds{1}\{1<s<k\}$, $k\geq 2$, and $\tilde{F}(s)=s^{-d}\mathds{1}\{s>1\}$. Let also $F_{k}(\mathbf{x})=\tilde{F}_{k}(|\mathbf{x}|)$ and $F(\mathbf{x})=\tilde{F}(|\mathbf{x}|)$. Note that $\{F_{k},k\geq 2\}$ is a sequence of functions such that
@@
$\sum_{k=1}^{d}z_{k}\frac{\partial\hat{f}(\mathbf{z})}{\partial z_{k}}=\theta\hat{f}(\mathbf{z})\hat{\nu}(\mathbf{z})-\theta\hat{f}(\mathbf{z}).$